OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, 1) and (4, 13)? y - 1 = -2(x - 4) y - 1 = 2(x + 2) y - 13 = -2(x + 2) y - 13 = 2(x + 4)

OpenStudy (anonymous):

DO YOU KNOW HOW TO FIND THE SLOPE?

OpenStudy (anonymous):

Sorry for the caps, I'm not that excited about rise-over-run

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

but i dont know how to find the answer with all the equations as answers :(

OpenStudy (anonymous):

Alright so first we need the slope, and once we do, we can get rid of some answers bc the slopes are different on some numbers compared to others

OpenStudy (anonymous):

Alright find the slope for me first and tell me what you get

OpenStudy (anonymous):

the slope is 2

OpenStudy (anonymous):

@doulikepiecauseidont

OpenStudy (anonymous):

Ok, so we're left with B and D since the A and C have slopes of -2

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

\[\Huge y-y_1=m(x-x_1)\] where m=slope (x1,y1) is a point

OpenStudy (anonymous):

If we try B we see that it works and if we try D we see that when we do the (x-x1) part we should get (x-4) but our answers gives us (x+4) if we had -4 as the point it would be (x-(-4)) which equals (x+4)

OpenStudy (anonymous):

oh so its B

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

THANK YOUUUUUU

OpenStudy (anonymous):

Yw